An isotope of carbon, 6C14, has a half-life of 5730 y. How much of a 4.0-g sample of the isotope will remain after 22,920 y?


1
Expert's answer
2021-03-14T19:21:41-0400

Let’s use the formula for radioactive decay:


"N=N_0e^{-\lambda t},""N=N_0e^{-\dfrac{0.693}{T_{1/2}} t},""N=4.0\ g\cdot e^{-\dfrac{0.693}{5730\ y}\cdot22920\ y}=0.25\ g."